# High-dimensional maximum-entropy phase space tomography using normalizing flows

### v1

@@ III Numerical experiments
The following numerical experiments demonstrate that MENT-Flow solutions approach MENT solutions in 2D phase space. Subsequent experiments demonstrate that MENT-Flow can fit complicated 6D phase space distributions to large measurement sets in reasonable time and that entropic regularization keeps the reconstruction close to the prior distribution. We tended to use ground-truth distributions without linear interplane correlations and simple transformations rather than realistic accelerator models. [^6] Equivalently, we assume we know the covariance matrix $\bm{\mathbf{\Sigma}}=\langle\bm{\mathbf{x}}\bm{\mathbf{x}}^{T}\rangle=\bm{\mathbf{V}}\bm{\mathbf{V}}^{T}$ and reconstruct the distribution in normalized coordinates $\bm{\mathbf{x}}_{n}=\bm{\mathbf{V}}^{-1}\bm{\mathbf{x}}$ by setting $\mathcal{M}_{k}\rightarrow\mathcal{M}_{k}\bm{\mathbf{V}}$. The covariance matrix is usually overdetermined by the measurements—for example, three measurements determine the $2\times 2$ covariance matrix—so that all distributions that fit the data share the same covariance matrix. In these cases it is reasonable to fit the covariance matrix first. We also chose to maximize the entropy relative to a Gaussian prior distribution ($\rho_{*}(\bm{\mathbf{x}})$) and used a Gaussian base distribution, making the maximum-entropy flow a linear transformation. [^7] A Gaussian prior may be a reasonable choice for accelerator applications: the prior has no interplane dependence and can expand to approximate a uniform distribution; (ii) any known elements of the $n\times n$ covariance matrix can be used to define the Gaussian prior; (iii) beams are typically clustered in phase space and approximately Gaussian at equilibrium; (iv) a Gaussian prior can be used to limit the beam size in dimensions that are weakly constrained by the data; and (v) the relative entropy has a maximum of zero. All training and analysis code is available in an online repository [27].
@@
Our first experiment tests the model performance in 2:1 phase space tomography. We assume a linear accelerator lattice, composed of drifts and quadrupole magnets, such that a symplectic transfer matrix $\bm{\mathbf{M}}$ approximates the dynamics. The transfer matrix can be decomposed as
@@
Fig. 2 shows reconstructions from a varying number of projections, comparing MENT, MENT-Flow, and an unregularized neural network (NN).
Figure 2: 2D reconstructions from varying numbers of evenly spaced 1D projections. The top rows plot samples from the true distribution, MENT reconstruction, MENT-Flow reconstruction, and unregularized neural network reconstruction. Faint lines show the evenly spaced projection angles, increasing from 1 on the left column to 7 on the right column. The projections of the four distributions are plotted on the horizontal axes on the bottom rows.
@@
Fig. 2 illustrates that entropy maximization is a conservative approach to the reconstruction problem. All reconstructed features are implied by the data. In contrast, the distributions in the bottom rows fit the data but are unnecessarily complex. Of course, reconstructions from one or two projections are bound to fail, no matter the algorithm, but these cases are still useful because they demonstrate MENT’s logical consistency: given only the marginal distributions and an uncorrelated prior distribution, MENT returns the product of the marginals. On the other extreme, with enough data, the feasible distributions will differ only in minor details. MENT shines in intermediate cases where the measurements contain just enough information to constrain the distribution’s primary features. For example, the continuous spiral structure develops rapidly with the number of views.
Fig. 2 also illustrates the flow’s capacity to represent complex distributions despite its restriction to invertible transformations. This examples focuses on spiral patterns, which are characteristic of nonlinear dynamics. (Additional examples are included in the supplemental material.) It is important to note that, while our analysis focuses on the beam core, low-density regions can also impact accelerator performance [28]. Flows can struggle to model distribution tails when trained via maximum-likelihood [29]. Our example distribution does not have significant halo and we do not report the agreement at this level, but preliminary studies indicate that the Kullback-Leibler (KL) divergence may enhance dynamic range relative to, i.e., the mean absolute error when fitting the data.
@@
It is more challenging to design and evaluate 6D numerical experiments. First, we do not know the information-maximizing set of 6D phase space transformations. In the 2D case, if the transformations are linear, there is a clear connection between a single transformation parameter (the phase advance) and the reconstruction quality. 4D or 6D symplectic transformations require many more parameters [30], and it is unclear which parameters are relevant. Note also that the ideal set of transformations depends on the distribution and on the number of measured dimensions. Here, we restrict our attention to 1D projections, i.e., 6:1 reconstructions, for which we can plausibly obtain the ideal set of linear phase space transformations by generalizing the notion of“projection angle”. A 1D projection axis is defined by a point on the unit $n$-sphere; if the distribution is spherically symmetric, we expect the optimal projection axes to be uniformly distributed on the sphere. [^8] It is unclear what “evenly spaced projection angles” correspond to when the integration planes have dimension $n-2$. In the limit of many such projections, the reconstruction will converge to the true distribution [31].
@@
Second, there are many possible 6D distributions, and establishing reconstruction accuracy requires high-dimensional visualization or statistical distance metrics. As described shortly, we selected distributions with clear high-dimensional structure and leveraged complete sets of pairwise projections and limited sets of partial projections, i.e., projections of slices, to facilitate the visualization task.
@@
Third, we cannot determine the distance from the reconstructed distribution to the true maximum-entropy distribution without an analytic solution. We point to Fig. 2 as evidence that the flow’s entropy estimate can approximate the true entropy, and we continue to train an unregularized neural network on the same data to show that additional solutions can exist far from our prior.
@@
Our first high-dimensional experiment, shown in Figs. 3-4, reconstructs a seven-mode Gaussian mixture distribution from 25, then 100 random 1D projections. This reconstruction used an approximately uniform prior and the same flow architecture as our 2D experiment.
@@
Figure 3: Reconstruction of 6D Gaussian mixture distribution from 25 random 1D projections. (a-b) Simulated/true projections. (c-d) 1D/2D projections of reconstructed 6D distribution. Reconstructed/true 2D projections are plotted on the lower/upper triangular; 1D projections are overlayed on the diagonal.
@@
Figure 4: Reconstruction of 6D Gaussian mixture distribution from 100 random 1D projections. (a-b) Simulated/true projections. (c-d) 1D/2D projections of reconstructed 6D distribution. Reconstructed/true 2D projections are plotted on the lower/upper triangular; 1D projections are overlayed on the diagonal.
We draw the following conclusions. (i) Flow-based models can represent complex 6D densities far from their unimodal base densities. All simulated measurements agree with the ground truth. Charged particle beams are often smooth and unimodal; therefore, we presume that flow-based models will be sufficient for most applications in accelerator physics. (ii) Some complex, high-dimensional distributions require large numbers of measurements to reconstruct with high precision. Of course, the required precision varies with application. In our case, the distribution’s primary features are largely determined from only 25 measurements in Fig. 3, but the entropy penalty eliminates unnecessary high-frequency terms in the distribution function. These terms are still present in the unregularized reconstruction when we quadruple the number of measurements in Fig. 4. (iii) MENT-Flow can scale to large measurement sets in reasonable time.
@@
The Gaussian mixture distribution has little overlap between modes, so mismatch between the true and reconstructed distribution is obvious from low-dimensional views. This is not always the case. Direct measurements of low-energy hadron beams have unveiled hollow structures in high-dimensional phase space which are not evident in low-dimensional views. [^9] Measurements at the Spallation Neutron Source (SNS) Beam Test Facility (BTF) [1] show space-charge-driven hollowing in 3D and 5D projections of the 6D phase space distribution. This motivates us to test tomographic reconstruction algorithms on distributions with hidden internal structure. To this end, an $n$-dimensional “rings” distribution serves as the ground truth in Fig. 5-6; particles populate two concentric $n$-spheres with radii $r_{2}=2r_{1}$, and the radii are perturbed with Gaussian noise to generate a smooth density.
Figure 5: Reconstruction of 6D “rings” distribution from 25 random 1D projections. (a-b) Simulated/true projections. (c-d) Reconstructed 2D projections within shrinking 4D ball of radius $\tilde{r}_{\perp}=\sqrt{x^{2}+x^{\prime 2}+y^{2}+y^{\prime 2}}$ in unplotted dimensions.
Figure 6: Reconstruction of 6D “rings” distribution from 100 random 1D projections. (a-b) Simulated/true projections. (c-d) Reconstructed 2D projections within shrinking 4D ball of radius $\tilde{r}_{\perp}=\sqrt{x^{2}+x^{\prime 2}+y^{2}+y^{\prime 2}}$ in unplotted dimensions.
@@
The entropy-regularized distribution maintains the spherical symmetry of the Gaussian prior, flattening and eventually inverting its radial density profile to fit the data. [^10] The one-dimensional projections in Figs. 5-6 are nearly Gaussian. Klartag [32] proved that almost all $m$-dimensional projections of an isotropic (no linear correlations) $n$-dimensional log-concave distribution function are nearly Gaussian when $n\gg m$. Many distributions commonly used in accelerator modeling are log-concave, such as the $n$-dimensional Gaussian, Waterbag (uniformly filled ball), and KV (uniformly filled sphere) distributions. A practical implication of this theorem is that small fluctuations in the $m$-dimensional projections have a greater impact on the $n$-dimensional reconstructed distribution as $|n-m|$ increases—for instance, completely inverting the density profile from peaked to hollow. Thus, we found that later training epochs can significantly change the distribution while only slightly decreasing the loss function. It follows that, for certain distributions, there may be some value of $|n-m|$ for which $n$:$m$ tomography is practically impossible. The sliced views reveal an internal structure—a dense core surrounded by a low-density cloud—that is better approximated by MENT-Flow when measurements are scarce. In addition to injecting unnecessary correlations between planes, the unregularized solution ejects all particles from the core. Surprisingly, adding additional measurements does not solve the problem and generates two distinct modes in the reconstructed density. Using a different random seed to define the measurement axes can generate different patterns, but the hollowing and splitting just described are typical. Note that this internal structure is not obvious from the full 2D projections in the left column of Figs. 5-6. In experimental settings, we often have strong a priori assumptions of smoothness and unimodality based on beam dynamics simulations and theory. By encoding this background knowledge in MENT-Flow’s prior distribution, we are likely to obtain more reliable high-dimensional reconstructions.
@@ IV Conclusion and extensions
MENT-Flow is a promising approach to high-dimensional maximum-entropy phase space tomography. Numerical experiments demonstrate consistency with known 2D maximum-entropy solutions and the ability to fit complex 6D phase space distributions to large measurement sets. Our numerical experiments also emphasize the importance of regularization and uncertainty quantification in high-dimensional phase space tomography. We demonstrated that large numbers of 1D projections are insufficient to reconstruct certain complex 6D distributions with high precision. Future work should apply the method to more realistic distributions and phase space transformations. Additionally, future work should study the importance of entropy regularization when the measurements are 2D projections (which we expect to contain exponentially more information per measurement) or come from alternative diagnostics [33]. At the same time, alternative maximum-entropy algorithms should be investigated to address a wider range of problems.
@@
MENT-Flow’s primary shortcomings are its slow sampling speed, its large number of parameters, and its restriction to differentiable particle dynamics. Sampling is approximately 50% slower than a conventional neural network, and the total runtime is increased by the need to solve multiple subproblems to approach the maximum-entropy distribution from below. This motivates the search for even more efficient flows and sample-based entropy estimates. Note, however, that the total training time is already reasonable, ranging from 3 to 15 minutes on a single GPU in our experiments (depending on the number of projections, phase space dimension, batch size, and penalty parameter updates).
@@
Flow-based phase space tomography may be able to approach maximum-entropy solutions to unexplored problems. First, MENT-Flow may be able to reconstruct $n$-dimensional distributions from high-dimensional projections ($m>2$). Such problems are of theoretical interest but also have some practical relevance. 3D and 4D projections can be measured relatively quickly using slit-screen-dipole measurement systems in low-energy hadron accelerators [1, 2, 3]. Projections may be partial, i.e., defined within a slice along other dimensions. The differentiable kernel density estimation needed to simulate the projections will perform poorly in these cases, so we propose to draw samples from the measured projections and minimize a statistical distance between these samples and samples from the flow.
Next, a particularly interesting application of maximum-entropy tomography is to intense hadron beams. Since space charge can generate highly nonlinear self-fields that depend on the 3D spatial distribution, knowledge of the full 6D phase space distribution is needed to predict the evolution of intense beams. A common measurement setup is to place the reconstruction location upstream of the measurement location, thereby allowing low-dimensional measurements under varying transport conditions (e.g. quadrupole scan) to obtain more information about the beam at the reconstruction point. Introducing transport renders the tomography problem a significant challenge for intense beams because the forward process depends on self-fields of the unknown phase space distribution. With MENT-Flow, it may be possible to include space charge in the beam dynamics using differentiable space charge models [34]. We also plan to pursue tomography under intense space charge with conventional MENT, which may provide a benchmark for MENT-Flow results.
Finally, quantifying reconstruction uncertainty will be a crucial step for the operational use of high-dimensional phase space tomography. Of particular interest are Bayesian methods, where a prior probability is assigned to every possible phase space distribution and updated after observing the data. (The principle of maximum entropy claims that the prior should increase as a function of entropy [13], so the strategy in this paper remains correct if we must select a single distribution.) Several authors have used generative models to solve inverse image reconstruction problems via Bayesian posterior sampling [35]. Perhaps similar methods could be applied to our case.

### v3

@@ III Numerical experiments
The following numerical experiments demonstrate that MENT-Flow solutions approach MENT solutions in 2D phase space. Subsequent experiments demonstrate that MENT-Flow can fit complicated 6D phase space distributions to large measurement sets in reasonable time and that entropic regularization keeps the reconstruction close to the prior. To simplify the examples, we focused on linear phase space transformations rather than more realistic accelerator models. We also tended to use ground-truth distributions without linear interplane correlations, highlighting nonlinear features. [^4] Equivalently, we assume we know the covariance matrix $\bm{\mathbf{\Sigma}}=\langle\bm{\mathbf{x}}\bm{\mathbf{x}}^{T}\rangle=\bm{\mathbf{V}}\bm{\mathbf{V}}^{T}$, where $\mathbf{V}$ is a symplectic matrix, and reconstruct the distribution in normalized coordinates $\bm{\mathbf{x}}_{n}=\bm{\mathbf{V}}^{-1}\bm{\mathbf{x}}$ by setting $\mathcal{M}_{k}\rightarrow\mathcal{M}_{k}\bm{\mathbf{V}}$. The covariance matrix is usually overdetermined by the measurements—for example, three measurements determine the $2\times 2$ covariance matrix—so that all distributions that fit the data share the same covariance matrix. In these cases, it is reasonable to fit the covariance matrix first. We chose to maximize the entropy relative to a Gaussian prior. [^5] A Gaussian prior may be a reasonable choice for accelerator applications: (i) the prior has no interplane dependence and can expand to approximate a uniform distribution; (ii) any known elements of the $n\times n$ covariance matrix can be used to define the Gaussian prior; (iii) beams are typically clustered in phase space and approximately Gaussian at equilibrium; (iv) a Gaussian prior can be used to limit the beam size in dimensions that are weakly constrained by the data. The flow’s base distribution is also a Gaussian, so the entropy penalty pushes the flow toward an identity or scaling transformation.
Our normalizing flow architecture is described in the previous section. The flow consists of five layers. Each layer is an autoregressive transformation, where the 1D transformation along each dimension is a rational-quadratic spline with 20 knots; the function values and derivatives at the knots are parameterized by a masked neural network with 3 hidden layers of 64 hidden units. Note that increasing the model size should not lead to overfitting since we train via maximum entropy, not maximum likelihood.
We compare MENT-Flow to MENT. Our MENT implementation uses linear interpolation to evaluate the Lagrange functions at any location on the projection axes, and we simulate the projections by numerical integration. We also compare to an unregularized neural network (NN) whose only aim is to fit the data. The NN is a standard fully connected feedforward network with 3 hidden layers of 32 hidden units and $\tanh$ activation functions.
We used $2\times 10^{4}$ samples to estimate the entropy and projections. During each epoch, we trained the normalizing flow using 400 iterations of the Adam optimizer. After each epoch, we multiplied the penalty parameter by $1.5$. We stopped training when
$\langle D\rangle=\frac{1}{K}\sum_{k=1}^{K}{D[{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}}),\tilde{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})]}<\epsilon,$ (21)
where $\langle D\rangle$ is the average divergence between the simulated and measured projections, $K$ is the number of measurements, and $\epsilon\approx 10^{-4}$ is a threshold. Other hyperparameter values are found in [31], which contains the code to reproduce the figures in this paper.
@@
Our first experiment tests the model performance in 2:1 phase space tomography. We assume an accelerator composed of drifts and quadrupole magnets, such that a symplectic transfer matrix $\bm{\mathbf{M}}$ approximates the dynamics. The transfer matrix can be decomposed as
@@
Fig. 2 shows reconstructions from a varying number of projections, comparing MENT, MENT-Flow, and the unregularized neural network (NN).
Figure 2: 2D reconstructions from evenly spaced 1D projections. The top four rows plot samples from the true distribution, MENT reconstruction, MENT-Flow reconstruction, and NN reconstruction. Faint lines show the evenly spaced projection angles, increasing from 1 in the left column to 7 in the right column. In the bottom rows, the distributions are projected onto the measurement axes. (The four profiles overlap in most cases.)
@@
Fig. 2 illustrates that entropy maximization is a conservative approach to the reconstruction problem. All reconstructed features are implied by the data. In contrast, the distributions in the bottom rows fit the data but are unnecessarily complex. Of course, reconstructions from one or two projections are bound to fail if the prior is uninformative, but these cases are still useful because they demonstrate MENT’s logical consistency: given only the marginal distributions and an uncorrelated prior, the posterior is the product of the marginals. On the other extreme, with enough data, the feasible distributions differ only in minor details. MENT shines in intermediate cases where the measurements contain just enough information to constrain the distribution’s primary features. For example, the continuous spiral structure develops rapidly with the number of views in Fig. 2.
Fig. 2 also illustrates the flow’s capacity to represent complicated distributions despite the restriction to invertible transformations. This example focuses on spiral patterns, which are characteristic of nonlinear dynamics. (Additional examples are included in the supplemental material.) It is important to note that, while our analysis focuses on the beam core, low-density regions can also impact accelerator performance [32]. Flows can struggle to model distribution tails [33]. Our ground-truth distribution does not have significant halo and we do not report the agreement at this level; however, preliminary studies indicate the Kullback-Leibler (KL) divergence may enhance dynamic range relative to, i.e., the mean absolute error when fitting data.
Fig. 3 plots the entropy and data mismatch terms during training.
Figure 3: Entropy and data mismatch during training.
The end of each epoch is clear from the sharp jumps in the loss curves when the penalty parameter $\mu$ increases. We aimed to keep the penalty parameter updates as small as possible. More aggressive update schedules did not lead to dramatically different results, but we did not explore this in detail. We did not see significant improvements using more sophisticated Augmented Lagrangian (AL) methods [17, 34]. A more important choice seems to be the stopping criteria, as increasing $\mu$ can eventually cause ill-conditioning. Our stopping condition ($\langle D\rangle\leq 10^{-4}$ in Eq. (21)) was chosen based on visual comparison of the simulated and true projections. We are not sure if the ideal stopping condition can be determined automatically.
@@
It is more difficult to design and evaluate high-dimensional numerical experiments. First, establishing reconstruction accuracy requires high-dimensional visualization or statistical distance metrics. We selected ground-truth distributions with clear high-dimensional structure and leveraged complete sets of pairwise projections and limited sets of partial projections (projections of slices) to aid the visualization.
@@
Second, we cannot determine the distance from the reconstructed distribution to the true maximum-entropy distribution without an analytic solution. We point to Fig. 2 as evidence that the entropy penalty can push the MENT-Flow solution close to the exact solution. We also continued to train an unregularized neural network on the same data to show that additional solutions can exist far from the prior.
Third, for a given beamline and a fixed number of measurements, we do not yet know how to find the information-maximizing set of 6D phase space transformations. In 2:1 tomography, if the transformations are linear, the reconstruction quality is tied to a single parameter (the projection angle). There is no such connection in $n$:2 tomography when $n>3$, as there is no obvious analog of the projection angle in these cases. Here, to demonstrate the method, we instead restrict our attention to 1D projections. A 1D projection axis can be specified by a point on the unit sphere; if the distribution is spherically symmetric, we hypothesize that the optimal projection axes are uniformly spaced on the sphere. In 2D, this leads to evenly spaced projection angles between 00 and $\pi$ radians. In our numerical experiments, we approximated this condition by randomly sampling points from a uniform distribution on the sphere. The points will not be uniformly spaced, but in the limit of many projections, the reconstruction should converge to the true distribution [35].
@@
Our first high-dimensional experiment, shown in Figs. 4-5, reconstructs a seven-mode Gaussian mixture distribution (a superposition of seven Gaussian distributions, each with a random mean and variance) from random 1D projections. Fig. 4 uses 25 projections and Fig. 5 uses 100 projections. This reconstruction used the same flow architecture as the 2D experiments. The NN architecture was changed to 2 layers of 50 units, still with $\tanh$ activation functions.
@@
Figure 4: Reconstruction of a 6D Gaussian mixture distribution from 25 random 1D projections. The MENT-Flow reconstruction on the left is compared to the NN reconstruction on the right. (a-b) Simulated projections (blue) vs. measured projections (red). (c-d) Low-dimensional views of the reconstructed distribution (blue) and the ground-truth distribution (red). 1D profiles are plotted on the diagonal subplots. 2D projections are plotted on the off-diagonal subplots.
@@
Figure 5: Reconstruction of a 6D Gaussian mixture distribution from 100 random 1D projections. The MENT-Flow reconstruction on the left is compared to the NN reconstruction on the right. (a-b) Simulated projections (blue) vs. measured projections (red). (c-d) Low-dimensional views of the reconstructed distribution (blue) and the ground-truth distribution (red). 1D profiles are plotted on the diagonal subplots. 2D projections are plotted on the off-diagonal subplots.
We draw the following conclusions. (i) Normalizing flows can represent complicated 6D distributions far from the unimodal base distribution. All simulated measurements match the training data. Charged particle beams are often smooth and unimodal, so this example represents a challenging case. Therefore, flow-based models are likely sufficient for many applications in accelerator physics. (ii) MENT-Flow can simultaneously fit a large number of measurements. (iii) The entropy penalty works as intended. The entropy-regularized solution fits the data just as well as the NN solution but eliminates high-frequency terms in the distribution function. The MENT-Flow solution is much closer to the smooth prior.
@@
The Gaussian mixture distribution has little overlap between modes, so mismatch between the true and reconstructed distribution is obvious from low-dimensional views. Hollow structures in high-dimensional phase space are not always evident from low-dimensional views. As an example, measurements at the Spallation Neutron Source (SNS) Beam Test Facility (BTF) show space-charge-driven hollowing in 3D and 5D projections of the 6D phase space distribution [1, 2, 3]. This motivates us to consider distributions with hidden internal structure. To this end, an $n$-dimensional “rings” distribution serves as the ground truth in Fig. 6-7; particles populate two concentric $n$-spheres with radii $r_{2}=2r_{1}$, and the radii are perturbed with Gaussian noise to generate a smooth density.
Figure 6: Reconstruction of 6D “rings” distribution from 25 random 1D projections. The MENT-Flow reconstruction on the left is compared to the NN reconstruction on the right. (a-b) Simulated projections (blue) vs. measured projections (red). (c-d) 2D projections of the 6D distribution $\rho(x,x^{\prime},y,y^{\prime},z,z^{\prime})$ within a shrinking 4D ball in the $x$-$x^{\prime}$-$y$-$y^{\prime}$ plane. We define a ball of radius $\tilde{r}_{\perp}$ by ${r}_{\perp}\leq\tilde{r}_{\perp}$, where ${r}_{\perp}=\sqrt{x^{2}+x^{\prime 2}+y^{2}+y^{\prime 2}}$. Therefore, we write the projected density as $\rho(z,z^{\prime}|r_{\perp}\leq\tilde{r}_{\perp})$. The ball shrinks from left to right. The largest radius (on the left) selects nearly all particles, while the smallest radius (on the right) selects particles near the core.
Figure 7: Reconstruction of 6D “rings” distribution from 100 random 1D projections. The MENT-Flow reconstruction on the left is compared to the NN reconstruction on the right. (a-b) Simulated projections (blue) vs. measured projections (red). (c-d) 2D projections of the 6D distribution $\rho(x,x^{\prime},y,y^{\prime},z,z^{\prime})$ within a shrinking 4D ball in the $x$-$x^{\prime}$-$y$-$y^{\prime}$ plane. We define a ball of radius $\tilde{r}_{\perp}$ by ${r}_{\perp}\leq\tilde{r}_{\perp}$, where ${r}_{\perp}=\sqrt{x^{2}+x^{\prime 2}+y^{2}+y^{\prime 2}}$. Therefore, we write the projected density as $\rho(z,z^{\prime}|r_{\perp}\leq\tilde{r}_{\perp})$. The ball shrinks from left to right. The largest radius (on the left) selects nearly all particles, while the smallest radius (on the right) selects particles near the core.
@@
The entropy-regularized solution maintains the spherical symmetry of the Gaussian prior, flattening and eventually inverting its radial density profile to fit the data. [^6] The one-dimensional projections in Figs. 6-7 are nearly Gaussian. Klartag [36] proved that almost all $m$-dimensional projections of an isotropic (no linear correlations) $n$-dimensional log-concave distribution function are nearly Gaussian when $n\gg m$. Many distributions commonly used in accelerator modeling are log-concave, such as the $n$-dimensional Gaussian, Waterbag (uniformly filled ball), and KV (uniformly filled sphere) distributions. A practical implication of this theorem is that small fluctuations in the $m$-dimensional projections have a greater impact on the $n$-dimensional reconstructed distribution as $n-m$ increases—for instance, completely inverting the density profile from peaked to hollow. Thus, we found that later training epochs can significantly change the distribution while only slightly decreasing the loss function. It follows that, for certain distributions, there may be some value of $n-m$ for which $n$:$m$ tomography is practically impossible. The sliced views reveal an internal structure—a dense core surrounded by a low-density cloud—that MENT-Flow better approximates when measurements are scarce. In addition to injecting unnecessary correlations between planes, the unregularized solution ejects all particles from the core. Surprisingly, adding additional measurements does not solve the problem and generates two distinct modes in the reconstructed density. Using a different random seed to define the measurement axes can generate different patterns, but the hollowing and splitting just described are typical. Note that this internal structure is not obvious from the full 2D projections in the left column of Figs. 6-7.
@@ IV Conclusion and extensions
In conclusion, MENT-Flow is a promising approach to high-dimensional phase space tomography. Numerical experiments demonstrate consistency with known 2D maximum-entropy solutions and the ability to fit complex 6D distributions to large measurement sets. In the 6D tests, although there are no available benchmarks, we found that entropic regularization pulls the solution closer to the prior. Thus, MENT-Flow is an effective way to incorporate prior information in high-dimensional reconstructions. Our numerical experiments also emphasize the potential importance of uncertainty quantification in high-dimensional tomography, as we found that some distributions can only be reconstructed from large numbers of 1D measurements. Future work should apply MENT-Flow to more realistic distributions, accelerator models, and diagnostics, especially 2D projections. Future work should also aim to extend MENT to higher dimensions to serve as a benchmark.
@@
MENT-Flow has several limitations. First, particle sampling is over 50% slower than a conventional neural network, and the total runtime is inflated by the need to solve multiple subproblems to approach the maximum-entropy distribution from below. This motivates the search for more efficient flows and sample-based entropy estimates. Note that our training times ranged from 5 to 20 minutes on a single GPU, depending on the number of projections, phase space dimension, batch size, and penalty parameter updates. Second, MENT-Flow maximizes the entropy using a penalty method that does not generate exact solutions and requires a hand-tuned penalty parameter schedule to avoid ill-conditioning. It is unclear whether this process can be automated or whether alternative strategies can better prevent ill-conditioning. Third, MENT-Flow does not attach uncertainty to its output.
@@
We now discuss possible extensions to new problems. First, it may be possible to fit $n$-dimensional distributions to $m$-dimensional projections when $m>2$. This problem is of theoretical interest but also has some practical relevance. 3D and 4D projections can be measured relatively quickly using slit-screen-dipole measurement systems in low-energy hadron accelerators [1, 2, 3]. We propose to draw samples from the measured projections and minimize a differentiable statistical distance between these samples and samples from the normalizing flow.
Second, an interesting application of maximum-entropy tomography is to intense hadron beams, in which particles respond to both applied and self-generated electromagnetic fields. Phase space tomography is a significant challenge for such beams because the forward process depends on the unknown initial distribution. We do not know if the maximum-entropy distribution is unique in this case. Including space charge in the GPSR forward process may be possible using differentiable space charge solvers [37].
Finally, quantifying reconstruction uncertainty is a crucial step for the operational use of phase space tomography. In this paper, we defined a prior probability distribution $\rho(\bm{\mathbf{x}})$ over the phase space coordinates $\bm{\mathbf{x}}$. We may also define a prior over the space of distribution functions. In practice, the phase space distribution is parameterized by $\bm{\mathbf{\theta}}\in\mathbb{R}^{N}$, so we may write the prior as $\mathcal{P}(\bm{\mathbf{\theta}})$. The set of discretized measurements, i.e., histograms, can be expressed as a another parameter vector $\bm{\mathbf{d}}\in\mathbb{R}^{M}$. Bayesian inference provides the update from prior to posterior:
$\mathcal{P}(\bm{\mathbf{\theta}}|\bm{\mathbf{d}})=\frac{\mathcal{P}(\bm{\mathbf{d}}|\bm{\mathbf{\theta}})\mathcal{P}(\bm{\mathbf{\theta}})}{\mathcal{P}(\bm{\mathbf{d}})},$ (25)
where $\mathcal{P}(\bm{\mathbf{d}}|\bm{\mathbf{\theta}})$ is the likelihood, encoding the forward model, and $\mathcal{P}(\bm{\mathbf{d}})$ is a normalizing constant. The principle of maximum entropy implies that we should prefer higher entropy phase space distributions. Gull and Skilling [14] addressed this problem for image reconstruction and argued that the prior $\mathcal{P}(\bm{\mathbf{\theta}})$ should be proportional to the exponential of the entropy $H(\bm{\mathbf{\theta}})$. Given this prior, the strategy in this paper finds the maximum, or mode, of the posterior $\mathcal{P}(\bm{\mathbf{\theta}}|\bm{\mathbf{d}})$. The full posterior maps the entire solution space, encoding the reconstruction uncertainty. Future work could attempt to sample from a Gaussian approximation of the posterior at its maximum [14]. Alternatively, it may be possible to sample from a more accurate approximation of the posterior using Markov Chain Monte Carlo (MCMC) or machine learning methods [38].